# Arrangements of circles supported by small chords and compatible with natural real algebraic functions

### v1

@@ Abstract.
A volume of a subset of the Euclidean space is defined naturally from the standard Euclidean metric (if it is a Lebesgue measurable set for example).
@@
For example, for circles in the plane we have abused the concept that they are sufficiently small with no explicit exposition: we can guess this naturally and we can formulate by using arguments via sufficiently small real number $\epsilon>0$. A CS-region (of $\mathcal{S}$) of a certain class is sufficiently small if the area of $\overline{D}$ and the length of $\overline{D}-D$, which are defined of course, are sufficiently small (we can formulate by using arguments via sufficiently small real numbers $\epsilon>0$ for areas and lengths). This is seen as a way of natural formulation.
We present lengthy and important arguments as Theorem 1. This is important in presenting Theorem 3 for example.
We also note that different from general articles and preprints on mathematics, we present most of proofs of our new results in our present paper associated with the statements. We can check most of our arguments by investigating carefully.
@@
For an NI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, choose a point $x_{j^{\prime},0}\in\overline{D_{\mathcal{S}}}$ contained in a single circle $S_{j}$ and choose a sufficiently small circle ${S^{\prime}}_{x_{j^{\prime},0}}$ centered there and bounds the closed disk ${D^{\prime}}_{x_{j^{\prime},0}}$. We have the unique chord of $\mathcal{S}$ connecting the two points $x_{j^{\prime},0,1}$ and $x_{j^{\prime},0,2}$ of ${S^{\prime}}_{x_{j^{\prime},0}}\bigcap S_{j}\bigcap\overline{D_{\mathcal{S}}}={S^{\prime}}_{x_{j^{\prime},0}}\bigcap S_{j}:=\{x_{j^{\prime},0,1},x_{j^{\prime},0,2}\}$. We have a chord of $\mathcal{S}$ which is also a subset of ${D^{\prime}}_{x_{j^{\prime},0}}$ and the two chord boundary points of which are denoted by $x_{j^{\prime},0,3}$ and $x_{j^{\prime},0,4}$, respectively. We can also have one such that the five points $x_{j^{\prime},0,1}$, $x_{j^{\prime},0,3}$, $x_{j^{\prime},0}$, $x_{j^{\prime},0,4}$ and $x_{j^{\prime},0,2}$ are mutually distinct and located on the smooth connected curve ${D^{\prime}}_{x_{j^{\prime},0}}\bigcap S_{j}$ in this order.
@@
    First, for an arbitrary chord of $\mathcal{S}$ obtained as before, either it is always a subset of $\overline{D_{\mathcal{S}}}\bigcap{D^{\prime}}_{x_{j^{\prime},0}}$ or a subset of $({\mathbb{R}}^{2}-D_{\mathcal{S}})\bigcap{D^{\prime}}_{x_{j^{\prime},0}}$. In the former (latter) case, we call a point $x_{j^{\prime},0}$ a convex (resp. concave) $D_{\mathcal{S}}$-point. We call such a chord of $\mathcal{S}$ a chord at $(D_{\mathcal{S}},x_{j^{\prime},0})$.
@@
    1. (a)
      Let a point $x_{j^{\prime},0}$ be not a horizontal pole or a vertical pole of $S_{j}$. Then for the chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ before, we consider the representation of the form $\{p_{1}+t(p_{2}-p_{1})\mid 0\leq t\leq 1\}$ and for $(p_{1,2,1},p_{1,2,2}):=p_{2}-p_{1}$, $p_{1,2,1}\neq 0$ and $p_{1,2,2}\neq 0$ must be satisfied and the signs of $p_{1,2,1}$ and $p_{1,2,2}$ are either of the following for any of these chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ and these presentations: the signs are always same or they are always different and this only depends on the point $x_{j^{\prime},0}$.
    2. (b)
      Let a point $x_{j^{\prime},0}$ be a horizontal pole of $S_{j}$. Then for the chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ before, we consider the representation of the form $\{p_{1}+t(p_{2}-p_{1})\mid 0\leq t\leq 1\}$ and for $(p_{1,2,1},p_{1,2,2}):=p_{2}-p_{1}$, $p_{1,2,1}\neq 0$ must be satisfied. Furthermore, we can have all cases of $p_{1,2,1}$ and $p_{1,2,2}$ of the following: the case $p_{1,2,2}\neq 0$ with the signs being same, the case $p_{1,2,2}\neq 0$ with the signs being different and the case $p_{1,2,2}=0$.
    3. (c)
      Let a point $x_{j^{\prime},0}$ be a vertical pole of $S_{j}$. Then for the chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ before, we consider the representation of the form $\{p_{1}+t(p_{2}-p_{1})\mid 0\leq t\leq 1\}$ and for $(p_{1,2,1},p_{1,2,2}):=p_{2}-p_{1}$, $p_{1,2,2}\neq 0$ must be satisfied. Furthermore, we can have all cases of $p_{1,2,1}$ and $p_{1,2,2}$ of the following: the case $p_{1,2,1}\neq 0$ with their signs being same, the case $p_{1,2,1}\neq 0$ with their signs being different and the case $p_{1,2,1}=0$.
  3. (3)
    We can have the unique CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ depending on $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ such that for the notation ${\bigcup}_{S_{j}\in J}C_{\mathcal{S},S_{j}}\bigcup{S_{\mathcal{S}}}^{\prime}$ (${S_{\mathcal{S}}}^{\prime}\subset S_{\mathcal{S}}$) before $C_{\mathcal{S},S_{j}}:=S_{\mathcal{S},x_{j^{\prime},0}}$, $J:=\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$, ${S_{\mathcal{S}}}^{\prime}:=C_{D_{\mathcal{S}},x_{j^{\prime},0}}$, and $S_{\mathcal{S}}:={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$. Here we can also have the unique CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ depending on $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ such that for the notation ${\bigcup}_{S_{j}\in J}C_{\mathcal{S},S_{j}}\bigcup{S_{\mathcal{S}}}^{\prime}$ before $C_{\mathcal{S},S_{j}}=S_{\mathcal{S},\overline{x_{j^{\prime},0}}}$, $J=\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$, ${S_{\mathcal{S}}}^{\prime}=C_{D_{\mathcal{S}},x_{j^{\prime},0}}$, and $S_{\mathcal{S}}={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$. We call the former resulting CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$. We call the latter resulting CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ an unsupported CS region at $(D_{\mathcal{S}},x_{j^{\prime},0})$.
    1. (a)
      Let $x_{j^{\prime},0}$ be a convex $D_{\mathcal{S}}$-point. If a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ is sufficiently small, then we can choose $S_{x_{j^{\prime}},r_{j^{\prime}}}:=S_{\mathcal{S}}$ and $E_{x_{j^{\prime}},r_{j^{\prime}}}:=D_{x_{j^{\prime}},r_{j^{\prime}}}$ in Definition 3 to have another NI arrangement ${\mathcal{S}}^{\prime}$ of circles. For an NCI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an NCI arrangement of circles. For an MBC arrangement $(\mathcal{S},D_{\mathcal{S}})$, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an MBC arrangement. If an unsupported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ is sufficiently small, then we can choose $S_{x_{j^{\prime}},r_{j^{\prime}}}:=S_{\mathcal{S}}$ and $E_{x_{j^{\prime}},r_{j^{\prime}}}:={\mathbb{R}}^{2}-D_{x_{j^{\prime}},r_{j^{\prime}}}$ in Definition 3 to have another NI arrangement ${\mathcal{S}}^{\prime}$ of circles. For an NCI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an NCI arrangement of circles. For an MBC arrangement $(\mathcal{S},D_{\mathcal{S}})$, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an MBC arrangement.
    2. (b)
@@
      Let $x_{j^{\prime},0}$ be a concave $D_{\mathcal{S}}$-point. We restrict the class of supported CS-regions at $(D_{\mathcal{S}},x_{j^{\prime},0})$ to that of CS-regions of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ containing the following uniquely defined bounded CS-region of $\mathcal{S}$: for the notation ${\bigcup}_{S_{j}\in J}C_{\mathcal{S},S_{j}}\bigcup{S_{\mathcal{S}}}^{\prime}$ before $C_{\mathcal{S},S_{j}}$ is a curve containing $x_{j^{\prime},0}$ with $J=\{S_{j}\}$, ${S_{\mathcal{S}}}^{\prime}=C_{D_{\mathcal{S}},x_{j^{\prime},0}}$, and $S_{\mathcal{S}}={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$. This subclass is not empty. If a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ of this new class is sufficiently small, then we can choose $S_{x_{j^{\prime}},r_{j^{\prime}}}:=S_{\mathcal{S}}={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$ and $E_{x_{j^{\prime}},r_{j^{\prime}}}:={\mathbb{R}}^{2}-D_{x_{j^{\prime}},r_{j^{\prime}}}$ in Definition 3 to have another NI arrangement ${\mathcal{S}}^{\prime}$ of circles. For an NCI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an NCI arrangement of circles. For an MBC arrangement $(\mathcal{S},D_{\mathcal{S}})$, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an MBC arrangement.
@@
As commented above, we can check this by investigating carefully. We only show explicit cases FIGUREs 1, 2, 3 and 4.
Figure 1. Chords of $\mathcal{S}$ of three types for Theorem 1 (2c).
Figure 2. A supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ for Theorem 1 (3a) and Theorem 2 (1).
Figure 3. An unsupported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ for Theorem 1 (3a) and Theorem 2 (2b).
Figure 4. The intersection of a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ and $\overline{D_{\mathcal{S}}}$ for Theorem 1 (3b) and Theorem 5.
@@
    In the case Theorem 1 (2a) or (2b), the change is explained as either of the following.
@@
    In the case Theorem 1 (2c) and (3a), ${\pi}_{2,1,1}(x_{j^{\prime},0})$ is the minimum or the maximum in the image of the function ${\pi}_{2,1,1}{\mid}_{\overline{D_{\mathcal{S}}}}$ and $x_{j^{\prime},0}$ is the unique point such that the value is the minimum or the maximum. Let ${\pi}_{2,1,1}(x_{j^{\prime},0})$ be the minimum in the image. The change is explained as either of the following and both of these two can be realized.
@@
      This is for the case of considering a sufficiently small CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ in Theorem 1 (3a) except an unsupported CS-region $D^{\prime}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ with a supporting secant of $(D^{\prime},\mathcal{S})$ being defined for the case $p_{1,2,1}=0$ in Theorem 1 (2c) with the unique extension of the chord of $\mathcal{S}$.
@@
      This is for the case of considering a sufficiently small unsupported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ in (3a) with a supporting secant of $(D^{\prime},\mathcal{S})$ being defined for the case $p_{1,2,1}=0$ in Theorem 1 (2c).
@@
As commented, we can prove Theorem 2 by investigating carefully one by one. This is also important in [7]. ∎
@@
For an NI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, choose a point $x_{j^{\prime},0}\in\overline{D_{\mathcal{S}}}$ contained in two distinct circles $S_{j_{1}},S_{j_{2}}\in\mathcal{S}$ and choose a sufficiently small circle ${S^{\prime}}_{x_{j^{\prime},0}}$ centered there and bounds the closed disk ${D^{\prime}}_{x_{j^{\prime},0}}$. We have the unique chord of $\mathcal{S}$ connecting the two points $x_{j^{\prime},0,1}$ and $x_{j^{\prime},0,2}$ of ${S^{\prime}}_{x_{j^{\prime},0}}\bigcap\overline{D_{\mathcal{S}}}\bigcap(S_{j_{1}}\bigcup S_{j_{2}}):=\{x_{j^{\prime},0,1},x_{j^{\prime},0,2}\}$. We have a chord of $\mathcal{S}$ which is also a subset of ${D^{\prime}}_{x_{j^{\prime},0}}$ and the two chord boundary points of which are denoted by $x_{j^{\prime},0,3}$ and $x_{j^{\prime},0,4}$, respectively. We also have one such that the five points $x_{j^{\prime},0,1}$, $x_{j^{\prime},0,3}$, $x_{j^{\prime},0}$, $x_{j^{\prime},0,4}$ and $x_{j^{\prime},0,2}$ are mutually distinct and located on the piecewise smooth connected curve ${D^{\prime}}_{x_{j^{\prime},0}}\bigcap(S_{j_{1}}\bigcup S_{j_{2}})$ in this order.
  1. (1)
    First, for an arbitrary chord of $\mathcal{S}$ as before, it is always a subset of $\overline{D_{\mathcal{S}}}\bigcap{D^{\prime}}_{x_{j^{\prime},0}}$. We call such a chord of $\mathcal{S}$ a chord at $(D_{\mathcal{S}},x_{j^{\prime},0})$.
  2. (2)
    For any circular secant $S_{\mathcal{S}}$ of $\mathcal{S}$ passing the two chord boundary points of any chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$, we have a CS-region $D$ of $\mathcal{S}$ such that $x_{j^{\prime},0}$ is contained in the closure $\overline{D}$, that $S_{\mathcal{S}}$ is a supporting secant of $(D,\mathcal{S})$ and that a labeling set for $(D,\mathcal{S})$ is a two-element set $\{S_{j_{1}},S_{j_{2}}\}$. For a point $(p_{j_{1},1},p_{j_{1},2})\neq x_{j^{\prime},0}$ in a supporting subset $C_{\mathcal{S},S_{j_{1}}}$ of $(D,S_{j_{1}},\mathcal{S})$ and a point $(p_{j_{2},1},p_{j_{2},2})\neq x_{j^{\prime},0}$ in a supporting subset $C_{\mathcal{S},S_{j_{2}}}$ of $(D,S_{j_{2}},\mathcal{S})$, we have a vector $((p_{j_{1},1},p_{j_{1},2})-x_{j^{\prime},0},(p_{j_{2},1},p_{j_{2},2})-x_{j^{\prime},0})\in{\mathbb{R}}^{4}$. Then we uniquely have the corresponding tuple of the signs of these components, whose values are not 00, depending on $(\mathcal{S},D_{\mathcal{S}})$, the point $x_{j^{\prime},0}$ and the ordered pair $(S_{j_{1}},S_{j_{2}})$ only, and independent of a supporting secant $S_{\mathcal{S}}$ of $(D,\mathcal{S})$, and a supporting subset $C_{\mathcal{S},S_{j_{i}}}$ of $(D,S_{j_{i}},\mathcal{S})$ ($i=1,2$) for example. We call the tuple of the $4$ signs the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$.

### v2

@@ Abstract.
The volume of a subset of the Euclidean space is defined naturally from the standard Euclidean metric (if it is a Lebesgue measurable set for example).
@@
For example, for circles in the plane we have abused the concept that they are sufficiently small with no explicit exposition: we can guess this naturally and we can formulate by using arguments via sufficiently small real number $\epsilon>0$. A CS-region (of $\mathcal{S}$) of a certain class is sufficiently small if the area of $\overline{D}$ and the length of $\overline{D}-D$, which are defined of course, are sufficiently small (we can formulate by using arguments via sufficiently small real numbers $\epsilon>0$ for areas of the regions and lengths of the curves). This is seen as a way of natural formulation.
For an NI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, choose a point $x_{j^{\prime},0}\in\overline{D_{\mathcal{S}}}$ contained in a single circle $S_{j}$ and choose a sufficiently small circle ${S^{\prime}}_{x_{j^{\prime},0}}$ centered there and bounds the closed disk ${D^{\prime}}_{x_{j^{\prime},0}}$. We have the unique chord of $\mathcal{S}$ connecting the two points $x_{j^{\prime},0,1}$ and $x_{j^{\prime},0,2}$ of ${S^{\prime}}_{x_{j^{\prime},0}}\bigcap S_{j}\bigcap\overline{D_{\mathcal{S}}}={S^{\prime}}_{x_{j^{\prime},0}}\bigcap S_{j}:=\{x_{j^{\prime},0,1},x_{j^{\prime},0,2}\}$. We have a chord of $\mathcal{S}$ which is also a subset of ${D^{\prime}}_{x_{j^{\prime},0}}$ and the two chord boundary points of which are denoted by $x_{j^{\prime},0,3}$ and $x_{j^{\prime},0,4}$, respectively. We can also have one such that the five points $x_{j^{\prime},0,1}$, $x_{j^{\prime},0,3}$, $x_{j^{\prime},0}$, $x_{j^{\prime},0,4}$ and $x_{j^{\prime},0,2}$ are mutually distinct and located on the connected curve ${D^{\prime}}_{x_{j^{\prime},0}}\bigcap S_{j}$ in this order. We can check the following easily.
Proposition 1.
For an arbitrary chord of $\mathcal{S}$ obtained above, either it is always a subset of $\overline{D_{\mathcal{S}}}\bigcap{D^{\prime}}_{x_{j^{\prime},0}}$ or a subset of $({\mathbb{R}}^{2}-D_{\mathcal{S}})\bigcap{D^{\prime}}_{x_{j^{\prime},0}}$.
Definition 6.
For Proposition 1, in the former (latter) case, we call a point $x_{j^{\prime},0}$ a convex (resp. concave) $D_{\mathcal{S}}$-point. We call such a chord of $\mathcal{S}$ a chord at $(D_{\mathcal{S}},x_{j^{\prime},0})$.
We can check the following easily.
Proposition 2.
  1. (1)
    Let a point $x_{j^{\prime},0}$ be not a horizontal pole or a vertical pole of $S_{j}$. For the chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ above, we consider the representation of the form $\{p_{1}+t(p_{2}-p_{1})\mid 0\leq t\leq 1\}$. For $(p_{1,2,1},p_{1,2,2}):=p_{2}-p_{1}$, $p_{1,2,1}\neq 0$ and $p_{1,2,2}\neq 0$ must be satisfied and the signs of $p_{1,2,1}$ and $p_{1,2,2}$ are either of the following for any of these chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ and these presentations: the signs are always same or they are always different and this only depends on the pair $(D_{\mathcal{S}},x_{j^{\prime},0})$.
  2. (2)
    Let a point $x_{j^{\prime},0}$ be a horizontal pole of $S_{j}$. For the chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ above, we consider the representation of the form $\{p_{1}+t(p_{2}-p_{1})\mid 0\leq t\leq 1\}$. For $(p_{1,2,1},p_{1,2,2}):=p_{2}-p_{1}$, $p_{1,2,1}\neq 0$ must be satisfied. Furthermore, we can have all cases of $p_{1,2,1}$ and $p_{1,2,2}$ of the following: the case $p_{1,2,2}\neq 0$ with the signs being same, the case $p_{1,2,2}\neq 0$ with the signs being different, and the case $p_{1,2,2}=0$.
  3. (3)
    Let a point $x_{j^{\prime},0}$ be a vertical pole of $S_{j}$. For the chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ above, we consider the representation of the form $\{p_{1}+t(p_{2}-p_{1})\mid 0\leq t\leq 1\}$. For $(p_{1,2,1},p_{1,2,2}):=p_{2}-p_{1}$, $p_{1,2,2}\neq 0$ must be satisfied. Furthermore, we can have all cases of $p_{1,2,1}$ and $p_{1,2,2}$ of the following: the case $p_{1,2,1}\neq 0$ with their signs being same, the case $p_{1,2,1}\neq 0$ with their signs being different, and the case $p_{1,2,1}=0$.
We show an example for Proposition 2 (3) as FIGURE 1, for example.
Figure 1. Chords of $\mathcal{S}$ of three types for Proposition 2 (3).
We can have the unique CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ depending on $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ such that for the notation ${\bigcup}_{S_{j}\in J}C_{\mathcal{S},S_{j}}\bigcup{S_{\mathcal{S}}}^{\prime}$ (${S_{\mathcal{S}}}^{\prime}\subset S_{\mathcal{S}}$) before $C_{\mathcal{S},S_{j}}:=S_{\mathcal{S},x_{j^{\prime},0}}$, $J:=\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$, ${S_{\mathcal{S}}}^{\prime}:=C_{D_{\mathcal{S}},x_{j^{\prime},0}}$, and $S_{\mathcal{S}}:={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$. Here we can also have the unique CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ depending on $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ such that for the notation ${\bigcup}_{S_{j}\in J}C_{\mathcal{S},S_{j}}\bigcup{S_{\mathcal{S}}}^{\prime}$ before $C_{\mathcal{S},S_{j}}=S_{\mathcal{S},\overline{x_{j^{\prime},0}}}$, $J=\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$, ${S_{\mathcal{S}}}^{\prime}=C_{D_{\mathcal{S}},x_{j^{\prime},0}}$, and $S_{\mathcal{S}}={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$. For this, Definition 7 introduces classes of CS regions of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$.
Definition 7.
We call the former resulting CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$. We call the latter resulting CS-region of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ an unsupported CS region at $(D_{\mathcal{S}},x_{j^{\prime},0})$.
@@
    Let $x_{j^{\prime},0}$ be a convex $D_{\mathcal{S}}$-point.
    If a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ is sufficiently small, then we can choose $S_{x_{j^{\prime}},r_{j^{\prime}}}:=S_{\mathcal{S}}$ and $E_{x_{j^{\prime}},r_{j^{\prime}}}:=D_{x_{j^{\prime}},r_{j^{\prime}}}$ in Definition 3 to have another NI arrangement ${\mathcal{S}}^{\prime}$ of circles. For an NCI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an NCI arrangement of circles, in this case. For an MBC arrangement $(\mathcal{S},D_{\mathcal{S}})$, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an MBC arrangement, in this case.
    If an unsupported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ is sufficiently small, then we can choose $S_{x_{j^{\prime}},r_{j^{\prime}}}:=S_{\mathcal{S}}$ and $E_{x_{j^{\prime}},r_{j^{\prime}}}:={\mathbb{R}}^{2}-D_{x_{j^{\prime}},r_{j^{\prime}}}$ in Definition 3 to have another NI arrangement ${\mathcal{S}}^{\prime}$ of circles. For an NCI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an NCI arrangement of circles, in this case. For an MBC arrangement $(\mathcal{S},D_{\mathcal{S}})$, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an MBC arrangement, in this case.
@@
    Let $x_{j^{\prime},0}$ be a concave $D_{\mathcal{S}}$-point. We restrict the class of supported CS-regions at $(D_{\mathcal{S}},x_{j^{\prime},0})$ to that of CS-regions of $\{S_{\mathcal{S},C_{D_{\mathcal{S}},x_{j^{\prime},0}}}\}$ containing the following uniquely defined bounded CS-region of $\mathcal{S}$: for the notation ${\bigcup}_{S_{j}\in J}C_{\mathcal{S},S_{j}}\bigcup{S_{\mathcal{S}}}^{\prime}$ before $C_{\mathcal{S},S_{j}}$ is a curve containing $x_{j^{\prime},0}$ with $J=\{S_{j}\}$, ${S_{\mathcal{S}}}^{\prime}=C_{D_{\mathcal{S}},x_{j^{\prime},0}}$, and $S_{\mathcal{S}}={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$. This subclass is not empty.
    If a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ of this new class is sufficiently small, then we can choose $S_{x_{j^{\prime}},r_{j^{\prime}}}:=S_{\mathcal{S}}={S}_{C_{D_{\mathcal{S}},x_{j^{\prime},0}}}$ and $E_{x_{j^{\prime}},r_{j^{\prime}}}:={\mathbb{R}}^{2}-D_{x_{j^{\prime}},r_{j^{\prime}}}$ in Definition 3 to have another NI arrangement ${\mathcal{S}}^{\prime}$ of circles. For an NCI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an NCI arrangement of circles, in this case. For an MBC arrangement $(\mathcal{S},D_{\mathcal{S}})$, $({\mathcal{S}}^{\prime},D_{{\mathcal{S}}^{\prime}})$ is also an MBC arrangement, in this case.
@@
We can check this by investigating carefully. We only show explicit cases FIGUREs 2, 3 and 4.
Figure 2. A supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ for Theorem 1 (1) and Theorem 2 (1).
Figure 3. An unsupported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ for Theorem 1 (1) and Theorem 2 (2b).
Figure 4. The intersection of a supported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ and $\overline{D_{\mathcal{S}}}$ for Theorem 2 (2) and Theorem 5.
@@
    In the case Proposition 2 (1) or (2), the change is explained as either of the following.
@@
    In the case Proposition 2 (3) and Theorem 1 (1), ${\pi}_{2,1,1}(x_{j^{\prime},0})$ is the minimum or the maximum in the image of the function ${\pi}_{2,1,1}{\mid}_{\overline{D_{\mathcal{S}}}}$ and $x_{j^{\prime},0}$ is the unique point such that the value is the minimum or the maximum. Let ${\pi}_{2,1,1}(x_{j^{\prime},0})$ be the minimum in the image. The change is explained as either of the following and both of these two can be realized.
@@
      This is for the case of considering a sufficiently small CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ in Theorem 1 (1) except an unsupported CS-region $D^{\prime}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$ with a supporting secant of $(D^{\prime},\mathcal{S})$ being defined for the case $p_{1,2,1}=0$ in Proposition 2 (3) and as the unique extension of the chord of $\mathcal{S}$.
@@
      This is for the case of considering a sufficiently small unsupported CS-region at $(D_{\mathcal{S}},x_{j^{\prime},0})$ in Theorem 1 (1) with a supporting secant of $(D^{\prime},\mathcal{S})$ being defined for the case $p_{1,2,1}=0$ in Proposition 2 (3) and as the unique extension of the chord of $\mathcal{S}$.
@@
We can prove Theorem 2 by investigating carefully one by one. This is also important in [7]. ∎
For an NI arrangement $(\mathcal{S},D_{\mathcal{S}})$ of circles, choose a point $x_{j^{\prime},0}\in\overline{D_{\mathcal{S}}}$ contained in two distinct circles $S_{j_{1}},S_{j_{2}}\in\mathcal{S}$ and choose a sufficiently small circle ${S^{\prime}}_{x_{j^{\prime},0}}$ centered there and bounds the closed disk ${D^{\prime}}_{x_{j^{\prime},0}}$. We have the unique chord of $\mathcal{S}$ connecting the two points $x_{j^{\prime},0,1}$ and $x_{j^{\prime},0,2}$ of ${S^{\prime}}_{x_{j^{\prime},0}}\bigcap\overline{D_{\mathcal{S}}}\bigcap(S_{j_{1}}\bigcup S_{j_{2}}):=\{x_{j^{\prime},0,1},x_{j^{\prime},0,2}\}$. We have a chord of $\mathcal{S}$ which is also a subset of ${D^{\prime}}_{x_{j^{\prime},0}}$ and the two chord boundary points of which are denoted by $x_{j^{\prime},0,3}$ and $x_{j^{\prime},0,4}$, respectively. We also have one such that the five points $x_{j^{\prime},0,1}$, $x_{j^{\prime},0,3}$, $x_{j^{\prime},0}$, $x_{j^{\prime},0,4}$ and $x_{j^{\prime},0,2}$ are mutually distinct and located on the connected curve ${D^{\prime}}_{x_{j^{\prime},0}}\bigcap(S_{j_{1}}\bigcup S_{j_{2}})$ in this order.
Proposition 3.
For an arbitrary chord of $\mathcal{S}$ above, it is always a subset of $\overline{D_{\mathcal{S}}}\bigcap{D^{\prime}}_{x_{j^{\prime},0}}$. We call such a chord of $\mathcal{S}$ a chord at $(D_{\mathcal{S}},x_{j^{\prime},0})$.
For any circular secant $S_{\mathcal{S}}$ of $\mathcal{S}$ passing the two chord boundary points of any chord $C_{D_{\mathcal{S}},x_{j^{\prime},0}}$ at $(D_{\mathcal{S}},x_{j^{\prime},0})$, we have a CS-region $D$ of $\mathcal{S}$ such that $x_{j^{\prime},0}$ is contained in the closure $\overline{D}$, that $S_{\mathcal{S}}$ is a supporting secant of $(D,\mathcal{S})$ and that a labeling set for $(D,\mathcal{S})$ is a two-element set $\{S_{j_{1}},S_{j_{2}}\}$. For a point $(p_{j_{1},1},p_{j_{1},2})\neq x_{j^{\prime},0}$ in a supporting subset $C_{\mathcal{S},S_{j_{1}}}$ of $(D,S_{j_{1}},\mathcal{S})$ and a point $(p_{j_{2},1},p_{j_{2},2})\neq x_{j^{\prime},0}$ in a supporting subset $C_{\mathcal{S},S_{j_{2}}}$ of $(D,S_{j_{2}},\mathcal{S})$, we have a vector $((p_{j_{1},1},p_{j_{1},2})-x_{j^{\prime},0},(p_{j_{2},1},p_{j_{2},2})-x_{j^{\prime},0})\in{\mathbb{R}}^{4}$. Then we uniquely have the corresponding tuple of the signs of these components, whose values are not 00, depending on $(\mathcal{S},D_{\mathcal{S}})$, the point $x_{j^{\prime},0}$ and the ordered pair $(S_{j_{1}},S_{j_{2}})$ only, and independent of a supporting secant $S_{\mathcal{S}}$ of $(D,\mathcal{S})$, and a supporting subset $C_{\mathcal{S},S_{j_{i}}}$ of $(D,S_{j_{i}},\mathcal{S})$ ($i=1,2$), for example.
Definition 8.
We call the tuple of the $4$ signs above the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$.